### v2

# Direct measurement of isotope shifts in the barium 6s2 1S0 - 5d6p 3D transition

Abstract
We report the first direct measurement of isotope shifts of the barium 6s2 1S0 - 5d6p 3D 413nm electric quadrupole transition, which is utilized for efficient barium ion trapping via photoionization using a single coherent light source. The measured isotope shifts relative to 138Ba are MHz, MHz, MHz, and MHz for isotopes with atomic number 137, 136, 135 and 134, respectively. We verified the measured isotopes with King plot analysis and compared the result with the formerly known shifts inferred from previous studies on neighboring transitions. The results could be used for efficient isotope selective loading of low-abundant barium ions, while careful suppression of line broadening is required for successful isotopic selectivity.
I Introduction
Barium ions are among the promising platforms for fundamental science and quantum information technology due to their comparatively long visible wavelength transitions and long-lived metastable D level[1, 2, 3, 4, 5]. Recently, odd isotopes of barium ions with non-zero nuclear spin have attracted interest owing to their rich electronic structure and exhibited unmatched state preparation and measurement (SPAM) error[6, 7], potential for extended error correction schemes[8], and high-dimensional Hilbert space to store quantum information[9]. However, scaling the odd-isotope barium ion trap necessitates high isotope-selective ion trapping since their natural abundance is relatively low compared to the most abundant barium-138 isotope.
Among the various ion-loading methods, the resonant photoionization technique has a clear advantage in isotope selectivity[10, 11, 12, 13]. In this process, atoms are resonantly excited to an intermediate state with high-isotope selectivity, and an additional laser removes a valence electron to ionize the atoms completely. So far, three intermediate states have been utilized for resonant photoionization of barium atoms. The 6s2 1S0 - 6s6p 1P transition ( = 553.5nm) has a strong transition strength for efficient ion loading [14, 15]. However, coherent radiation on the wavelength is only produced by a frequency-doubled or dye laser, and the isotope shifts are not sufficiently large for high isotope selectivity [16]. The 6s6p 3P1 state has large hyperfine splittings that allow efficient isotope selective loading[17, 18], but the second excitation wavelength falls into the UV region, which reportedly causes photoelectric ionization in an isotope-insensitive manner. The 5d6p 3D intermediate state has high electronic energy, enabling two-photon ionization by a single coherent 413nm laser source[15, 19]. However, isotope shifts of 413nm transitions and isotope-selective loading using this transition have not yet been reported.
Herein, we report the first direct measurement of the isotope shifts in the 6s2 1S0 - 5d6p 3D 413nm E2 transition of atomic barium. We measured the laser-induced fluorescence spectrum excited by a resonant 413nm laser source, and the isotope shifts of five naturally abundant barium isotopes were extracted from the spectrum. The measured shifts are compared with the indirect result of previous spectroscopic studies on neighboring transitions. Finally, we discuss the feasibility of isotope-selective loading using the 5d6p 3D intermediate level based on observations.
II Experimental setup
Figure 1 shows the experimental schematics and related electronic structure of barium. The barium atomic beam was produced by resistively heating the Ba metal in a thermal oven, which is placed inside a UHV vacuum chamber with the pressure below Torr. The 413nm laser perpendicular to the atomic beam excites the barium atoms to 5d6p 3D state, and the excited atoms emit fluorescence photons by spontaneous decay. Atomic fluorescence is collected and imaged by a CMOS image sensor (IMX249, Sony) to investigate the excitation spectrum, while the linear sensor responsivity was verified separately (See Appendix B). The atomic beam from the oven diverges as it propagates, and the transverse velocity distribution of the region of interest depends on the size of the oven aperture and the distance from the oven nozzle. To reduce the velocity distribution along the laser beam direction at the interrogation location, we clamped the barium oven aperture and aligned its short axis with a diameter of 180m parallel to the laser beam. The distance from the oven nozzle to the excitation laser beam is 8.4mm, and the angular diameter of the clamped nozzle from the region excited by the laser is approximately 2.1 mrad.
The 413nm excitation laser light is generated by an external cavity diode laser (Toptical DL pro). The laser frequency was monitored by a high-precision Fizeau-type wavelength meter (HighFinesse WS-8) and stabilized within 1 MHz around the target frequency controlled by homemade Python software[20]. The wavelength meter was regularly calibrated with a reference 780nm laser locked to the rubidium-87 D2 transition to compensate for the wavelength reading drift. The resonant 413nm light excited the ground-state barium atoms to 5d6p 3D state with a lifetime of 17.40.5 ns, and the predominant decay channels of the excited state are 6s5d 3D1 (64%) and 3D2 (32%) levels, emitting 660nm and 667nm photons respectively. An optical 660nm bandpass filter with a bandwidth of 10nm before the image sensor blocks ambient lights, and the measured fluorescence is predominantly 660nm light emitted by atoms.
III Result & Discussion
Figure 2 (a) shows a selection of the atomic fluorescence images obtained. As the laser frequency varies, the spatial distribution of atomic fluorescence changes due to the Doppler shift along the laser propagation direction, and the horizontal distribution width becomes the minimum when the light is resonant to the atomic transition. By comparing the Doppler-broadened spatial distribution of the images, the resonance frequency of the most abundant 138Ba was found to be 725.258895 THz, and the interrogation area where the atomic beam and the excitation laser are perpendiculars was also marked (black square in Fig. 2 (a)) (See Appendix A). The laser-induced fluorescence was quantified by summing the fluorescence counts in the region of interest. Figure 2 (b) shows the excitation spectrum of the 6s2 1S0 - 5d6p 3D transition constructed from the fluorescence images collected with different laser frequencies. Each data point was obtained by averaging ten images, and the step size of the laser frequency was 5 MHz.
In theory, odd isotopes (137 and 135) (I=3/2) have three hyperfine sublevels in the excited 5d6p 3D state, whereas even isotopes (138, 136, and 134) have no hyperfine splitting. To label each fluorescence peak with the corresponding isotopes and hyperfine levels, we deployed the hyperfine splittings of the 5d6p 3D state reported in Ref. [21]. The hyperfine energies can be derived from the Hamiltonian
where is Planck’s constant, is the magnetic dipole interaction constant, is the electric quadrupole constant, is the total angular momentum, and is the nuclear spin. Ref. [21] reported and coefficients of 5d6p 3D as = 155.01.5 MHz, = -2.42.0 MHz, = 172.90.3 MHz, and = -3.50.6 MHz. By utilizing the derived splitting values, six peaks in the spectrum were successfully matched with their respective isotopes and hyperfine levels, as depicted in Fig. 2 (b).
To determine the precise isotope shifts, we fit the spectrum to the following fit model:
with the individual peak function , the linear combination of Voigt distribution and Gaussian distribution accounting the main atomic beam and background ambient atoms respectively. Here, and are the natural abundance and isotope shift of barium isotope (), is the hyperfine shift of the corresponding hyperfine number , and are the Lorentzian and Gaussian width of Voigt profile, and are Gaussian width and relative signal strength of ambient background barium atoms, and is the amplitude of even and odd isotope peaks due to different polarization response.
| Fit parameter | Fit value |
| 392.5 0.6 MHz | |
| 177.8 0.4 MHz | |
| 400.6 1.0 MHz | |
| 125.9 1.4 MHz | |
| 11.2 0.4 MHz | |
| 3.1 0.5 MHz | |
| 69.8 2.4 MHz | |
| 1.090.02 | |
| 0.33 0.02 |
The model agreed well with the spectrum data and the hyperfine shifts from Ref. [21], and the isotope shifts of 6s2 1S0 - 5d6p 3D were successfully extracted (Fig 2 and Table 1). We found the Lorentzian full-width of fluorescence peaks (22.40.8 MHz) are relatively large compared to the expected 5d6p 3D decay rate 9.15 MHz, while the extracted isotope shifts are consistent regardless of peak model (See Appendix C and D).
While this work reports the first direct observation of isotope shifts in the 6s2 1S0 - 5d6p 3D transition, the isotope shifts of neighboring transitions reported in previous studies may indirectly provide the shift values. Figure 3 and Table 2 summarize the neighboring transitions and their reported isotope shifts. The 6s2 1S0 - 6s6p 1P 553.5nm transition has been investigated in multiple studies [22, 23, 24], and the reported isotope shifts were consistent with each other. There were a few experimental studies and a theoretical study on transitions of 6s5d metastable states [25, 21, 26, 27] and by using reported isotope shifts of the transition of 6s5d 3D1 - 6s6p 1P1 and 6s5d 3D1 - 5d6p 3D , the 6s2 1S0 - 5d6p 3D isotope shifts may be inferred as Table 2. It is noteworthy that the isotope shifts of 6s5d 3D1 - 6s6p 1P 1108nm from the experimental work [25] and the theory work [27] differ from each other and hence inferred 6s2 1S0 - 5d6p 3D isotope shifts vary as well. We found that the even isotope (134 and 136) cases are relatively consistent with each other, exhibiting a mismatch within 15MHz. However, the odd isotope (135 and 137) results showed a discrepancy, especially with the experimental work of Dammalapati et al. [25].
| (A=138) (MHz) | |||||
| 138 | 137 | 136 | 135 | 134 | |
| 6s2 1S0 - 6s6p 1P1 (553.5nm) | |||||
| Bekk et al. [23] | 0.0 | 215.00.7 | 128.90.5 | 260.90.7 | 143.00.5 |
| Wijngaarden et al. [24] | 0.0 | 215.20.2 | 128.00.4 | 259.30.2 | |
| 6s5d 3D1 - 6s6p 1P1 (1107.6nm) | |||||
| Dammalapti et al. [25] | 0.0 | -1144 | 59.30.6 | -394 | 144.11.0 |
| Nazé et al.∗ [27] | 0.0 | 67.2 | 79.5 | 132.1 | 139.5 |
| 5d6p 3D1 - 6s5d 3D1 (659.5nm) | |||||
| Grundevik et al. [21] | 0.0 | 189.21.0 | 112.91.0 | 228.31.0 | 125.02.0 |
| 6s2 1S0 - 5d6p 3D (413.2nm) | |||||
| This work | 0.0 | 392.50.6 | 177.80.4 | 400.61.0 | 125.91.4 |
| Based on Dammalapti et al.[25] | 0.0 | 518.24.1 | 182.51.3 | 528.24.2 | 123.92.3 |
| Based on Nazé et al.∗ [27] | 0.0 | 337.01.2 | 162.31.1 | 357.11.2 | 128.52.1 |
| Barium ion (singly charged), 6s 2S1/2 - 6p 2P1/2 (493.4nm) | |||||
| Imgram et al. [28] | 0.0 | 272.20.2 | 180.20.2 | 350.00.2 | 223.20.2 |
We compared our result with previous studies using the King plot [29]. When we introduce the modified isotope shift and plot those shifts in different transitions for all possible pairs , the data points are expected to fall on a straight line, which is called the King plot. Figure 4 shows the King plot of isotope shifts of our result and previous studies listed in Table 2 with reference data of the 6s 2S1/2 - 6p 2P1/2 barium ion transition [28]. Our experimental data show clear King plot linearity with a slope of 1.63, and other 553.5nm and 659.5nm transitions also follow a linear relationship. For the 1107.8nm transition, the theoretical work [27] shows linearity, but the work by Dammalapati et al. [25] shows a mismatch from the expected linearity, as already noted in the study. We believe that our study may give a hint for a better understanding of the spectroscopic characteristics of barium atoms.
Lastly, Table 3 shows the 413nm spectrum peaks sorted by their center frequencies, which could be utilized for the resonant isotope-selective photoionization process. For odd isotopes (135 and 137), the F=3/2 state among three hyperfine states has less proximity to its neighboring peaks, making it a preferred choice for photoionization. The difference between the peaks of 23.9 MHz is slightly greater than the natural line broadening of 9.15 MHz. It could be possible to achieve efficient isotope-selective photoionization by carefully reducing the effects of Doppler broadening and other line broadening factors.
| Peak center (MHz) | Peak label | Peak separation (MHz) |
| -44.1 | 137, F=1/2 | |
| 44.1 | ||
| 0.0 | 138 | |
| 9.2 | ||
| 9.2 | 135, F=1/2 | |
| 117.3 | ||
| 126.5 | 134 | |
| 51.4 | ||
| 177.9 | 136 | |
| 45.3 | ||
| 223.2 | 137, F=3/2 | |
| 23.9 | ||
| 247.1 | 135, F=3/2 | |
| 384.5 | ||
| 631.6 | 135, F=5/2 | |
| 19.5 | ||
| 651.1 | 137, F=5/2 |
In conclusion, the isotope shifts of the barium 6s2 1S0 - 5d6p 3D transition were directly measured from laser-induced fluorescence spectroscopy of the barium atomic beam. We confirmed the validity of our measurement with King plot analysis and compared the measured isotope shifts with the previous studies on adjacent transition lines. All nine transition frequencies for different isotopes and hyperfine levels were resolved, and the result suggests careful parameter control is required for isotope-selective loading of low-abundant isotope using 413nm transition.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (2022R1C1C100375812, 2022M3H3A106307412, RS-2023-00302576) and Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT) (2022-0-01040).Appendix A Selection of the region of interest in the images
The atomic beam from the oven diverges as it propagates, and the angle between atomic velocity and laser beam varies by interrogation areas. To achieve a Doppler-free spectrum, we set the image interrogation area such that the atoms traverse perpendicular to the laser beam. First, we compared the transverse profiles of the acquired fluorescence images along the laser propagation direction. Due to the high abundance of barium-138, the fluorescence distribution is mainly represented by barium-138 fluorescence. We observed that when the laser is off-resonant to barium-138, the peak of the transverse profile appears where Doppler shift compensates the detuning between the laser and atomic transition and its width is broader due to Doppler shift by velocity distribution. Among the images with different laser frequencies, we found the image with the minimum spatial width and the corresponding frequency was MHz, the resonance of barium-138 transition. Also, we set the region of interest around the fluorescence peak of the resonant image, the square region of 10 10 pixels.
Appendix B Linearity check of the CMOS image sensor responsivity
To verify the linearity of the CMOS sensor responsivity, we observed the sensor response to 650nm laser diode light while varying light intensities. The image sensor settings, including gain and exposure time, were maintained at the same values as those used in the atomic fluorescence experiment, and the laser intensity was reduced to match the atomic fluorescence intensity by employing an ND filter with an optical density of 3.8. The data on Fig. 6 indicated linear sensor response with the coefficent of determination () for the linear model () of 0.997.
Appendix C The linewidths of the fluorescence spectrum
To estimate the effect of transverse velocity distribution on the spectrum linewidth, we first measured the velocity distribution of the atomic beam. We measured the atomic fluorescence spectrum at a point 2.5mm shifted from the normal incident point along the excitation laser direction, where the incident angle between the excitation laser and atoms was 0.35 rad. By analyzing the Doppler shifts of the spectrum and given geometry between the laser beam and atomic beam, the fit of spectral data to the Boltzmann distribution yielded the root-mean-squared atomic velocity of 4352 m/s and temperature of 10508 K (Fig. 7). From the derived velocity distribution, the transverse velocity profile of a collimated effusive atomic beam could be derived from the model below [30]
where, is an incomplete gamma function, is the oven aperture size, is the detection area size, , and with the distance between the oven and detection area. Figure 8 shows the transverse velocity distribution derived from the velocity distribution, setup geometry, and corresponding Doppler shifts. The velocity distribution’s full width at half maximum (FWHM) is 6.150.03 m/s, equivalent to Doppler shifts of 14.80.1 MHz.
The spectra with different excitation laser intensities were obtained to confirm the absence of power broadening, and no increase in spectral width was observed (Fig. 9). In addition, the interaction time between the laser field and atoms was about 3 s, and the corresponding transit time broadening was negligible.
Appendix D Comparison of different peak models
Table 4 shows the fit results with various fit models. Although the detailed peak shape differs by the models, the extracted isotope shift values agree within 1 MHz and the choice of peak model does not affect the main result of this work. With pure Voigt peak model (), the fitted peak is mostly Lorentzian, while the uncertainty of Gaussian width is large. We could fit the data well with Voigt with background Gaussian model () while the Lorentizan width is fixed to the natural linewidth ( MHz), although the mean squared error (MSE) was slightly higher than the unfixed case. The fitted Gaussian FWHM width is MHz, comparable with expected Doppler broadening calculated in Appendix C
| Fit parameters | ( fixed) | ||
| (MHz) | 391.3 1.3 | 392.5 0.6 | 392.5 0.8 |
| (MHz) | 176.7 0.9 | 177.8 0.4 | 178.2 0.5 |
| (MHz) | 399.2 2.3 | 400.6 1.0 | 401.2 1.3 |
| (MHz) | 118.6 3.0 | 125.9 1.4 | 126.6 1.4 |
| (MHz) | 14.2 0.3 | 11.2 0.4 | 4.58 |
| (MHz) | 0.1 21.0 | 3.1 0.5 | 9.0 0.2 |
| (MHz) | N/A | 69.8 2.4 | 60.4 1.3 |
| N/A | 0.33 0.02 | 0.68 0.02 | |
| MSE | 1.03 | 2.31 | 6.16 |