### v3

# Direct measurement of isotope shifts in the barium 6s2 1S0 - 5d6p 3D transition

Abstract
We report the direct measurement of isotope shifts of the barium 6s2 1S0 –5d6p 3D 413-nm electric quadrupole transition, which is utilized for efficient barium ion trapping via photoionization using a single coherent light source. The measured isotope shifts relative to 138Ba are MHz, MHz, MHz, and MHz for isotopes with atomic numbers 137, 136, 135, and 134, respectively. We verify the measured isotopes with King plot analysis and compare the result with the formerly known shifts inferred from previous studies on neighboring transitions. The results can be used for efficient isotope selective loading of low-abundant barium ions, while careful suppression of line broadening is required for successful isotopic selectivity.
I Introduction
Barium ions are among the promising platforms for fundamental science and quantum information technology due to their comparatively long visible wavelength transitions and long-lived metastable D level [1, 2, 3, 4, 5]. Recently, odd isotopes of barium ions with non-zero nuclear spin have attracted interest owing to their rich electronic structure and exhibited unmatched state preparation and measurement (SPAM) error [6, 7], potential for extended error correction schemes [8], and high-dimensional Hilbert space to store quantum information [9]. However, scaling the odd-isotope barium ion trap necessitates high isotope-selective ion trapping since their natural abundance is relatively low compared to the most abundant barium-138 isotope.
Among the various ion-loading methods, the resonant photoionization technique has a clear advantage in isotope selectivity [10, 11, 12, 13]. In this process, atoms are resonantly excited to an intermediate state with high-isotope selectivity, and an additional laser removes a valence electron to ionize the atoms completely. So far, three intermediate states have been utilized for resonant photoionization of barium atoms. The 6s2 1S0 –6s6p 1P transition ( = 553.5 nm) has a strong transition strength for efficient ion loading [14, 15]. However, coherent radiation on the wavelength is only produced by a frequency-doubled laser or dye laser, and the isotope shifts are not sufficiently large for high isotope selectivity [16]. The 6s6p 3P1 state has large hyperfine splittings that allow efficient isotope selective loading[17, 18], but the second excitation wavelength falls into the UV region, which reportedly causes photoelectric ionization in an isotope-insensitive manner. The 5d6p 3D intermediate state has high electronic energy, enabling two-photon ionization by a single coherent 413-nm laser source[15, 19]. However, isotope shifts of 413-nm transitions and isotope-selective loading using this transition have not yet been reported.
Herein, we report the direct measurement of the isotope shifts in the 6s2 1S0 –5d6p 3D 413-nm E2 transition of atomic barium. We measured the laser-induced fluorescence spectrum excited by a resonant 413-nm laser source, and the isotope shifts of five naturally abundant barium isotopes were extracted from the spectrum. The measured shifts are compared with the indirect result of previous spectroscopic studies on neighboring transitions. Finally, we discuss the feasibility of isotope-selective loading using the 5d6p 3D intermediate level based on observations.
II Experimental setup
Figure 1 shows the experimental schematics and related electronic structure of barium. The barium atomic beam was produced by resistively heating the Ba metal in a thermal oven, which is placed inside a UHV vacuum chamber with the pressure below Torr. The 413-nm laser perpendicular to the atomic beam excites the barium atoms to 5d6p 3D state, and the excited atoms emit fluorescence photons by spontaneous decay. The laser beam has a 0.6-mm beam diameter and 60 W of beam power and the polarization is perpendicular to the atomic beam. Atomic fluorescence is collected and imaged by a complementary metal oxide semiconductor (CMOS) image sensor (IMX249, Sony) to investigate the excitation spectrum, while the linear sensor responsivity was verified separately (see Appendix A). The atomic beam from the oven diverges as it propagates, and the transverse velocity distribution of the region of interest depends on the size of the oven aperture and the distance from the oven nozzle. To reduce the velocity distribution along the laser beam direction at the interrogation location, we clamped the barium oven aperture and aligned its short axis with a diameter of 180 m parallel to the laser beam. The distance from the oven nozzle to the excitation laser beam is 8.4 mm, and the angular diameter of the clamped nozzle from the region excited by the laser is approximately 2.1 mrad.
The 413-nm excitation laser light is generated by an external cavity diode laser (Toptical DL pro). The laser frequency was monitored by a high-precision Fizeau-type wavelength meter (HighFinesse WS-8) and stabilized near the target frequency controlled by homemade software [20]. While the laser is locked, the frequency error recorded by the wavelength meter has a standard deviation of 0.4 MHz, the same as the wavelength meter resolution. The wavelength meter was regularly calibrated with a reference 780-nm laser locked to the rubidium-87 D2 transition to compensate for the wavelength reading drift. The resonant 413-nm light excited the ground-state barium atoms to 5d6p 3D state with a lifetime of 17.40.5 ns, and the predominant decay channels of the excited state are 6s5d 3D1 (64%) and 3D2 (32%) levels, emitting 660-nm and 667-nm photons, respectively. An optical 660-nm bandpass filter with a bandwidth of 10 nm before the image sensor blocks ambient lights, and the measured fluorescence is predominantly 660-nm light emitted by atoms.
To acquire the atomic fluorescence spectrum, we scanned the laser frequency by sequentially updating the lock target of wavelength-meter-based frequency stabilization. Once the laser frequency was stabilized to each target, atomic fluorescence images were achieved from the CMOS sensor. The atomic oven was kept on during the process, and the atomic flux from the oven remained constant throughout the experiment. The entire data acquisition process was automated via network remote procedure calls (RPC) to minimize the system drift.
III Result and Discussion
Figure 2 (a) shows a selection of the atomic fluorescence images obtained. As the laser frequency varies, the spatial distribution of atomic fluorescence changes due to the Doppler shift along the laser propagation direction, and the horizontal distribution width becomes the minimum when the light is resonant to the atomic transition. By comparing the Doppler-broadened spatial distribution of the images, the resonance frequency of the most abundant 138Ba was found to be 725.258895(5) THz, and the interrogation area where the atomic beam and the excitation laser are perpendicular is also marked [black square in Fig. 2 (a)] (see Appendix B). The laser-induced fluorescence was quantified by summing the fluorescence counts in the region of interest. Figure 2 (b) shows the excitation spectrum of the 6s2 1S0 –5d6p 3D transition constructed from fluorescence images collected with different laser frequencies. Each data point is obtained by averaging ten images, and the step size of the laser frequency was 5 MHz.
In theory, odd isotopes (137 and 135) (I=3/2) have three hyperfine sublevels in the excited 5d6p 3D state, whereas even isotopes (138, 136, and 134) have no hyperfine splitting. To label each fluorescence peak with the corresponding isotopes and hyperfine levels, we deploy the hyperfine splittings of the 5d6p 3D state reported in Ref. [21]. The hyperfine energies can be derived from the Hamiltonian
where is Planck’s constant, is the magnetic dipole interaction constant, is the electric quadrupole constant, is the total angular momentum, and is the nuclear spin. The authors of Ref. [21] reported and coefficients of 5d6p 3D as = 155.01.5 MHz, = -2.42.0 MHz, = 172.90.3 MHz, and = -3.50.6 MHz. By utilizing the derived splitting values, six peaks in the spectrum were successfully matched with their respective isotopes and hyperfine levels, as depicted in Fig. 2 (b).
To determine the precise isotope shifts, we fit the spectrum to the following fit model (, is the barium-138 resonance frequency):
with the individual peak function , the linear combination of Lorentzian distribution and Gaussian distribution accounting for the main atomic beam and background ambient atoms, respectively. Here, and are the natural abundance and isotope shift of the barium isotope (), is the hyperfine shift of the corresponding hyperfine number , and are the Lorentzian and Gaussian widths of the Voigt profile, and are Gaussian widths and relative signal strength of ambient background barium atoms, and is the amplitude of even and odd isotope peaks due to different polarization response.
| Fit parameter | Fit value | Fit uncertainty | Total uncertainty |
| (MHz) | 392.9 | 0.6 | 0.9 |
| (MHz) | 178.1 | 0.4 | 0.8 |
| (MHz) | 401.4 | 1.0 | 1.2 |
| (MHz) | 125.3 | 1.2 | 1.3 |
| (MHz) | 725258894.7 | 0.2 | - |
| (MHz) | 12.3 | 0.2 | - |
| (MHz) | 73.3 | 2.5 | - |
| 0.30 | 0.02 | - |
The model agrees well with the spectrum data and the hyperfine shifts from Ref. [21], and the isotope shifts of 6s2 1S0 –5d6p 3D were successfully extracted (Fig. 2 and Table 1).
Other than uncertainty from the fit, our data have uncertainty originated from the frequency measurement via wavelength meter, which is separately measured to be 0.7 MHz (see Appendix C). Since we are mainly interested in the relative frequency difference between isotope peaks, any common mode errors as laser linewidth and possible angle between atom and laser beam have a negligible effect on the analysis. By combining the fit uncertainty and measurement uncertainty, we estimate the total uncertainty of our isotope shift measurement. We find the Lorentzian full-width of fluorescence peaks ( MHz) are relatively large compared to the expected 5d6p 3D decay rate 9.15 MHz, while the extracted isotope shifts are consistent regardless of peak model (see Appendixes D and E).
While this work reports the direct observation of isotope shifts in the 6s2 1S0 –5d6p 3D transition, the isotope shifts of neighboring transitions reported in previous studies may indirectly provide the shift values. Figure 3 and Table 2 summarize the neighboring transitions and their reported isotope shifts. The 6s2 1S0 –6s6p 1P 553.5-nm transition were investigated in multiple studies [22, 23, 24], and the reported isotope shifts were consistent with each other. There were a few experimental studies and a theoretical study on transitions of 6s5d metastable states [25, 21, 26, 27] and by using reported isotope shifts of the transition of 6s5d 3D1–6s6p 1P1 and 6s5d 3D1–5d6p 3D , the 6s2 1S0 –5d6p 3D isotope shifts may be inferred as Table 2. It is noteworthy that the isotope shifts of 6s5d 3D1–6s6p 1P 1108 nm from the experimental work [25] and the theory work [27] differ from each other and hence infer 6s2 1S0 –5d6p 3D isotope shifts vary as well. We find that the even isotope (134 and 136) cases are relatively consistent with each other, exhibiting a mismatch within 15 MHz. However, the odd isotope (135 and 137) results show a discrepancy, especially with the experimental work of Dammalapati et al. [25].
| (A=138) (MHz) | |||||
| 138 | 137 | 136 | 135 | 134 | |
| 6s2 1S0 –6s6p 1P1 (553.5 nm) | |||||
| Bekk et al. [23] | 0.0 | 215.00.7 | 128.90.5 | 260.90.7 | 143.00.5 |
| Wijngaarden et al. [24] | 0.0 | 215.20.2 | 128.00.4 | 259.30.2 | |
| 6s5d 3D1–6s6p 1P1 (1107.6 nm) | |||||
| Dammalapti et al. [25] | 0.0 | -1144 | 59.30.6 | -394 | 144.11.0 |
| Nazé et al.∗ [27] | 0.0 | 67.2 | 79.5 | 132.1 | 139.5 |
| 5d6p 3D1–6s5d 3D1 (659.5 nm) | |||||
| Grundevik et al. [21] | 0.0 | 189.21.0 | 112.91.0 | 228.31.0 | 125.02.0 |
| 6s2 1S0 –5d6p 3D (413.2 nm) | |||||
| This work | 0.0 | 392.90.9 | 178.10.8 | 401.41.2 | 124.31.3 |
| Based on Dammalapti et al.[25] | 0.0 | 518.24.1 | 182.51.3 | 528.24.2 | 123.92.3 |
| Based on Nazé et al.a [27] | 0.0 | 337.01.2 | 162.31.1 | 357.11.2 | 128.52.1 |
| Barium ion (singly charged), 6s 2S1/2–6p 2P1/2 (493.4nm) | |||||
| Imgram et al. [28] | 0.0 | 272.20.2 | 180.20.2 | 350.00.2 | 223.20.2 |
We compare our result with previous studies using the King plot [29]. When we introduce the modified isotope shift and plot those shifts in different transitions for all possible pairs , the data points are expected to fall on a straight line, which is called the King plot. Figure 4 shows the King plot of isotope shifts of our result and previous studies listed in Table 2 with reference data of the 6s 2S1/2–6p 2P1/2 barium ion transition [28]. Our experimental data show clear King plot linearity with a slope of 1.63, and other 553.5 nm and 659.5 nm transitions also follow a linear relationship. For the 1107.8-nm transition, the theoretical work [27] shows linearity, but the work by Dammalapati et al. [25] shows a mismatch from the expected linearity, as already noted in the study. We believe that our study may give a hint for a better understanding of the spectroscopic characteristics of barium atoms.
Lastly, Table 3 shows the 413nm spectrum peaks sorted by their center frequencies, which could be utilized for the resonant isotope-selective photoionization process. For odd isotopes (135 and 137), the F=3/2 state among three hyperfine states has less proximity to its neighboring peaks, making it a preferred choice for photoionization. The difference between the peaks of 23.9 MHz is slightly greater than the natural line broadening of 9.15 MHz. It could be possible to achieve efficient isotope-selective photoionization by carefully reducing the effects of Doppler broadening and other line broadening factors.
| Peak center (MHz) | Peak label | Peak separation (MHz) |
| –43.7 | 137, F=1/2 | |
| 43.7 | ||
| 0.0 | 138 | |
| 10.9 | ||
| 10.9 | 135, F=1/2 | |
| 113.3 | ||
| 124.2 | 134 | |
| 53.9 | ||
| 178.1 | 136 | |
| 45.4 | ||
| 223.5 | 137, F=3/2 | |
| 25.3 | ||
| 248.8 | 135, F=3/2 | |
| 384.5 | ||
| 633.3 | 135, F=5/2 | |
| 18.1 | ||
| 651.4 | 137, F=5/2 |
In conclusion, the isotope shifts of the barium 6s2 1S0 –5d6p 3D transition were directly measured from laser-induced fluorescence spectroscopy of the barium atomic beam. We confirmed the validity of our measurement with King plot analysis and compared the measured isotope shifts with the previous studies on adjacent transition lines. All nine transition frequencies for different isotopes and hyperfine levels were resolved, and the result suggests careful parameter control is required for isotope-selective loading of low-abundant isotope using 413nm transition.
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (Grants No. 2022R1C1C100375812, No. 2022M3H3A106307412, and No. RS-2023-00302576) and Institute of Information and Communications Technology Planning and Evaluation (IITP) grant funded by the Korean government(MSIT) (Grant No. 2022-0-01040).Appendix A Linearity check of the CMOS image sensor responsivity
To verify the linearity of the CMOS sensor responsivity, we observed the sensor response to 650nm laser diode light while varying light intensities. The image sensor settings, including gain and exposure time, were maintained at the same values as those used in the atomic fluorescence experiment, and the laser intensity was reduced to match the atomic fluorescence intensity by employing a neutral-density (ND) filter with an optical density of 3.8. The data on Fig. 5 indicated linear sensor response with the coefficient of determination () for the linear model () of 0.997.
Appendix B Selection of the region of interest in the images
The atomic beam from the oven diverges as it propagates, and the angle between atomic velocity and laser beam varies by interrogation areas. To achieve a Doppler-free spectrum, we set the image interrogation area such that the atoms traverse perpendicular to the laser beam. First, we compare the transverse profiles of the acquired fluorescence images along the laser propagation direction. Due to the high abundance of barium-138, the fluorescence distribution is mainly represented by barium-138 fluorescence. We observe that when the laser is off-resonant to barium-138, the peak of the transverse profile appears where Doppler shift compensates the detuning between the laser and atomic transition and its width is broader due to Doppler shift by velocity distribution. Among the images with different laser frequencies, we fin the image with the minimum spatial width and the corresponding frequency is THz, suggesting the resonance of barium-138 transition is THz (see Fig. 6). Also, we set the region of interest around the fluorescence peak of the resonant image, the square region of 10 10 pixels.
Appendix C Relative uncertainty of the wavelength meter measurement
To investigate the relative measurement uncertainty of the wavelength meter, we employ the reference 780-nm laser locked to rubidium D2 lines. The laser is stabilized by saturated-absorption spectroscopy signal with rubidium vapor cell, and the laser frequency can be locked to various hyperfine states of the 2P3/2 excited state. The transition frequencies are well known from other studies [30, 31, 32, 33, 34]. We measure the laser frequency with the wavelength meter while changing the lock target transition. Table 4 shows the laser frequency measurements for various target transitions. We change the lock set points to six different hyperfine levels of 85Rb, 87Rb P3/2 excited states and compare the measured laser frequency with the references. The measurement result shows consistent agreement within 2 MHz between the measured frequency and the known transition frequency. We measure the average offset of 0.8 MHz and the standard deviation of data discrepancy is 0.7 MHz while the spanned frequency range is larger than 1.3GHz. In our main experiment, we mainly focus on the isotope shifts which are frequency differences between two data points and hence common mode error can be ignored. Therefore, we conclude our relative uncertainty of wavelength meter measurement to be 0.7MHz.
| Isotope and | |||
| ground state | Excited state | (MHz) | (MHz) |
| 87Rb, | –78.5 | –79.1(2) | |
| 0.0 | –0.4(2) | ||
| 133.3 | 132.9(2) | ||
| 85Rb, | 1167.8 | 1165.9(1) | |
| 1199.5 | 1199.5(2) | ||
| 1259.8 | 1258.1(1) |
Appendix D linewidths of the fluorescence spectrum
To estimate the effect of transverse velocity distribution on the spectrum linewidth, we first measure the velocity distribution of the atomic beam. We measure the atomic fluorescence spectrum at a point 2.5mm shifted from the normal incident point along the excitation laser direction, where the incident angle between the excitation laser and atoms is 0.35 rad. By analyzing the Doppler shifts of the spectrum and given geometry between the laser beam and atomic beam, the fit of spectral data to the Boltzmann distribution yielded the root-mean-squared atomic velocity of 4352 m/s and temperature of 10508 K (Fig. 7). From the derived velocity distribution, the transverse velocity profile of a collimated effusive atomic beam could be derived from the model below [35]
where, is an incomplete gamma function, is the oven aperture size, is the detection area size, = kT/m, and with the distance between the oven and detection area. Figure 8 shows the transverse velocity distribution derived from the velocity distribution, setup geometry, and corresponding Doppler shifts. The velocity distribution’s full width at half maximum (FWHM) is 6.150.03 m/s, equivalent to Doppler shifts of 14.80.1 MHz.
The spectra with different excitation laser intensities are obtained to confirm the absence of power broadening, and no increase in spectral width is observed (Fig. 9). In addition, the interaction time between the laser field and atoms was about 3 s, and the corresponding transit time broadening is negligible.
Appendix E Comparison of different peak models
Table 5 shows the fit results with various fit function models. Among the various models, Lorentzian peak with Gaussian background () shows the lowest reduced of 1.56, suggesting the model agrees the most with the data. Although the detailed peak shape differs by the models, the extracted isotope shift values agree within 1 MHz and the choice of peak model does not affect the main result of this work. With the pure Voigt peak model (), the fitted peak is mostly Lorentzian, while the uncertainty of the Gaussian width is large. We could also fit the data with the Voigt with the background Gaussian model () while the Lorentizan width is fixed to the natural linewidth ( MHz), although the mean squared error (MSE) is slightly higher than the unfixed case. The fitted Gaussian FWHM width is MHz, comparable with the expected Doppler broadening calculated in Appendix D.
| Fit parameters | ( fixed) | |||
| (MHz) | 392.9 0.6 | 391.9 1.2 | 392.9 0.6 | 392.7 0.8 |
| (MHz) | 178.1 0.4 | 176.3 0.9 | 178.1 0.4 | 178.3 0.5 |
| (MHz) | 401.4 1.0 | 401.3 2.2 | 401.4 1.0 | 401.4 1.4 |
| (MHz) | 124.3 1.2 | 122.3 2.3 | 125.3 1.2 | 126.7 1.4 |
| (MHz) | 12.3 0.2 | 15.5 0.5 | 12.3 0.6 | 4.58 (fixed) |
| (MHz) | N/A | 0.1 42.6 | 0.1 29.2 | 9.0 0.2 |
| (MHz) | 73.3 2.5 | N/A | 73.3 2.7 | 60.3 1.3 |
| 0.30 0.02 | N/A | 0.30 0.03 | 0.68 0.02 | |
| MSE | 3.61 | 1.35 | 3.6 | 6.06 |
| Reduced | 1.56 | 4.69 | 1.57 | 2.45 |